# Perceptions of Moderators as a Large-Scale Measure of Online Community Governance

### v1

@@ Abstract
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet it is challenging to determine what governance strategies are most successful, as measuring the ‘success’ of governance is complex and nuanced. Furthermore, the incredible diversity in community topic, size, and membership all but guarantee that there is no ‘one-size-fits-all’ solution for community governance. In this work, we measure governance by assessing how community members publicly discuss their own moderators. We quantify perceptions of moderators through 1.89 million labeled posts and comments made on reddit over an 18 month period, and relate these perceptions to characteristics of community governance and to different actions that community moderators can take. We identify key differences between different types of communities, and highlight promising strategies for moderator teams. Amongst other findings, we show that positive perceptions of moderators are associated with other measures of community health, and that strict rule enforcement is perceived more favorably for certain topics, such as news communities, than others. We investigate what kinds of moderators have the most positive impact on the community when they join the mod team, and find that moderators who are active community members before and during their mod tenures result in the largest improvement of community members’ perceptions of moderators. We make all our models, datasets, and code public.
1 Introduction
@@
Millions of online communities rely on volunteer moderators to carry out governance. These moderators perform essential services that substantially shape their communities, including setting and enforcing rules (Fiesler et al. 2018; Chandrasekharan et al. 2018; Seering et al. 2019), soliciting feedback from and communicating their actions to the community (Jhaver, Bruckman, and Gilbert 2019; Zhang, Hugh, and Bernstein 2020), and recruiting new moderators to keep pace with community growth or to replace moderators who have stepped down. Some form of governance provided by moderators seems essential to the healthy functioning of almost all online communities (Kraut and Resnick 2012; Matias 2019a).
It is challenging to determine what styles of governance are most effective at ensuring high quality outcomes for online communities. The solution space for online governance is extraordinarily large, with moderator teams needing to make countless decisions about how to govern: how many moderators do they need, and how should they select them? What rules should they enforce, and how should they enforce them? What even is the community’s best interest, anyways? Compounding this challenge is the enormous scale and diversity of online communities, millions of which exist on many platforms, in many sizes, and with many topics and goals. Prior work has established that there is no ‘one-size-fits-all’ solution for improving online governance (Weld, Zhang, and Althoff 2024, 2022).
@@
This scale and diversity also presents an opportunity. If we can develop an effective method to scalably assess the success of community governance practices, we can leverage the natural diversity of these practices across thousands of communities in order to identify the most promising strategies for moderators. However, measuring the success of governance practices at scale is extremely difficult. Survey methods can be effective but expensive to deploy widely, challenging to run longitudinally, and impossible to ‘backfill’ historic data. Notions of ‘success’ are multifaceted, and while automated classifiers can be used to detect specific harms such as misinformation (Weld, Glenski, and Althoff 2021) or specific aspects of governance such as rules enforcement (Chandrasekharan et al. 2018; Fiesler et al. 2018), current methods are unable to quantify broader notions of successful governance.
@@
In this work, we measure online community governance by examining how community members themselves perceive their moderators. We develop a method to classify how community members discuss their moderators, publicly, within their communities (§3). We use this method to gather and characterize community members’ perceptions of moderators at a massive scale, enabling the largest study-to-date of governance practices that we are aware of. We label 1.89 million posts and comments from 8,477 unique subreddits across an 18-month period from January 2021-June 2021, and relate these data to different kinds of online communities, and to different actions that community mod teams can take.
@@
Our findings have important implications for moderator teams (§5). We show that communities with fewer than 5 daily posts+comments per moderator use 2.5×2.5$\times$2.5 × as much positive language to describe their mods as communities with 20×20$\times$20 × more posts+comments per mod. Yet our findings do not suggest that more strict moderation improves perception of moderators; in fact, for most types of communities, a 3 percentage point increase in removed content is associated with a 33% increase in negative language used to describe mods, although news-sharing communities are a notable exception to this trend (§5.1). We examine the impact of adding different types of moderators to a community, and find that moderators who are also active community members before and during their tenure as moderators are most strongly associated with improved perceptions of mods (§5.2-9).
@@
We discuss the limitations of our methods, and identify key areas for future work (§6). We make our models, datasets, and code public [^1] These materials are being prepared for publication. For immediate use, please contact the first author. to enable further research on this important topic.
2 Related Work
@@
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera, Halfaker, and Terveen 2009) and some gaming servers (Frey and Sumner 2019). On reddit and most popular social media platforms, however access is difficult to obtain, and, and requires obtaining authorization separately from the mods of every community to be studied (Li, Hecht, and Chancellor 2022a, b). In contrast, our method works for every community whose content is public.
@@
Because of these challenges, many researchers turn to surveys to study governance, sometimes qualitatively (Matias 2019a), and sometimes at a larger scale, for instance to quantify harassment of mods (Almerekhi, Kwak, and Jansen 2020) or moderator recruiting practices (Seering and Kairam 2022).
@@
Governance Through the Lens of Rules. Partially because of their more-public nature, rules are often used as a lens to study governance. Large-scale analyses of rules have been used to characterize governance at the platform level on reddit (Reddy and Chandrasekharan 2023; Fiesler et al. 2018) and on Wikipedia (Hwang and Shaw 2022), however it is challenging to infer much about governance in specific communities, as rules and their enforcement constitute only a small portion of governance activity. Surveys of community members’ attitudes towards rules have been used to understand members’ attitudes towards governance more broadly (Koshy et al. 2023), however this method is challenging to scale beyond a single community. Our method can easily be applied to thousands of communities at once.
@@
Moderation Strategies on Reddit. Rules and their enforcement can also be evaluated in terms of their impact on the community as a whole (Jhaver, Bruckman, and Gilbert 2019; Srinivasan et al. 2019; Jhaver, Rathi, and Saha 2023) and their embodiment of communitywide norms (Chandrasekharan et al. 2018). Beyond rules and their enforcement, researchers have also studied other moderator strategies on reddit, including platform level decisions such as community bans and quarantines (Chandrasekharan et al. 2020; Ribeiro et al. 2020), and community level actions such as user bans (Thomas et al. 2021) and stickied posts (Matias 2019b). Our work quantifies many of these strategies, including content removal, mod interactions with the community, and associates them with positive and negative moderation discourse.
@@
Measuring Outcomes in Online Communities. Researchers have studied the range of values that community members hold for their communities (Weld, Zhang, and Althoff 2024), yet it is challenging to accurately predict these values automatically (Weld, Zhang, and Althoff 2022). Large-scale embeddings can be used to understand community culture (Waller and Anderson 2020), yet not governance directly. Longitudinal work has examined how communities fare with massive growth (Lin et al. 2017) and the lifecycles of their members(Danescu-Niculescu-Mizil et al. 2013), while some research focuses on smaller scale outcomes at the conversation level (Bao et al. 2021). Our work complements this literature with a large scale method for studying governance-specific outcomes.
3 Methods
@@
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al. 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, and moderators. These attributes, along with the the fact that almost all content on reddit is publicly available, make it an ideal platform for studying community governance at a large scale.
@@
In this work we make extensive use of publicly available reddit data downloaded from the Pushshift dataset (Baumgartner et al. 2020). From these data, we detect and classify and detect posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). Additionally, we quantify moderators’ activity on the platform before and during their moderator tenures (such as the number of times a mod posted in a community before being appointed a moderator there, as well as activity by non-moderators in thousands of communities).
3.1 Computing Moderator Timelines
@@
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of each subreddit’s moderators using the Wayback Machine, a web archiving service provided by the Internet Archive. [^2] https://archive.org/ We start by scraping current moderator info pages for the 10,000 largest subreddits on the platform, directly from reddit. Each such page contains a list of the current moderators, in order of seniority, along with the exact timestamp that each moderator was added as a moderator to the subreddit. We then use the publicly accessible API from the Internet Archive’s Wayback Machine to scrape every archived copy of every subreddits’ moderator info page going back to 2010. We scrape 30,302 historical archived copies of moderator info pages, which, combined with the 10,000 present-day copies (one per subreddit), give us detailed timelines of who moderated each subreddit when, and when they started as a moderator. Due to the functionality of the Wayback Machine, we have higher temporal resolution (more archived snapshots) for larger and more popular subreddits, yet, because each moderator info page encodes the exact timestamp for each moderator’s start date, having fewer snapshots results only decreased accuracy of when moderators resign their posts, not when they are appointed — which is the primary focus of this paper. Since the exact end time of a moderator’s tenure can only be inferred from examining when they were removed from the list of moderators for each subreddit, we adopt a conservative strategy which deliberately underestimates the length of moderators’ tenures: we consider the end date of their tenure as that of the last snapshot for which they were still listed as a moderator. Our method only misses moderators whose appointment, entire tenure, and resignation all occur within the ‘gap’ between snapshots [^3] We believe this is rare, as the average mod tenure we can detect is longer than four years, and the mean gap between snapshots is 56 days.. After a final data-cleaning step to remove known bots, the resulting dataset of moderator timelines, which we make public, consists of 89,237 moderator tenures across 9,502 subreddits.
@@
Table 1: Our Classification Step model, a LLaMA 2 model fine tuned with QLoRA (Touvron et al. 2023; Dettmers et al. 2023), exceeds the performance of a retrieval-based few-shot classifier using GPT-4. Our model, which we make public, is alsomore scalable (as it can be deployed locally), more affordable, and more reliable (as it is not subject to prompt filtering), than GPT-4. This table compares $F_{1}$ scores for expert human labelers, retrieval-based few-shot GPT-4, and our model, alongside an empirical class distribution random baseline and a VADER-based classifier (Hutto and Gilbert 2014).
3.2 Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the 10,000 largest subreddits, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld, Zhang, and Althoff 2022). We used the manually-labeled dataset of 123 subreddits from Weld, Zhang, and Althoff (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld, Zhang, and Althoff 2022). We average across all responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
3.3 Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify posts and comments made by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects posts and comments which discuss moderators (differentiating them from those which discuss other uses of ‘mod(s)’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of these posts and comments into positive, neutral, negative, and exclude classes. We only include comments made by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not on how mods discuss themselves in public. Our pipeline, which we make public, was applied to the entire reddit dataset of public posts and comments made from January 2020 to June 2021.
Figure 1: Determining the sentiment with regards to the moderators of comments can be very challenging.
Detection Step Details (Step 2). To differentiate posts and comments which discuss moderators from those which use ‘mod(s),’ we fine-tuned a RoBERTa-based binary classifier (Liu et al. 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the entire reddit dataset and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88, and, when applied to the entire reddit dataset, left us with 330,356 posts and 2,426,049 comments which discuss moderators.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa. To manually label posts and comments for the sentiment classification task, two annotators worked together to iteratively refine a codebook (Appendix A), then independently labeled 200 randomly sampled posts and comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts and comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts and comments for use as a training set, of which 484 were randomly sampled, and additional 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa based model trained for the classification task.
@@ 3.3 Detecting and Classifying Community Members’ Perceptions of Moderators
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the appropriate moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators (e.g., discussion of the /r/cats moderators taking place on /r/cats). For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with posts or comments that erroneously slipped through the detection step. We trained on this ‘exclude’ class in addition to the classes used for downstream analyses: positive, neutral, and negative [^4] Additionally, we attempted to classify specific complaints about moderation: excessive moderation, insufficient moderation, and biased moderation, but found classifier performance not to be amenable for high-confidence downstream analyses..
To further enhance the performance of our model, we perform data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al. 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al. 2023) fine-tuned with QLoRA (Dettmers et al. 2023). LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step leave us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make this dataset public.
Accurately classifying the sentiment with regards to the moderators of a post or comment is an extremely challenging task that is often heavily reliant on context, background knowledge of the community in which the discussion is occurring, and requiring inside knowledge about the community’s norms and culture (Figure 1). As such, off-the-shelf sentiment classifiers, such as VADER, perform very poorly (Table 1). Even general performance LLMs such as GPT-4, which obtain state-of-the-art results on standard sentiment analysis benchmarks (Kheiri and Karimi 2023), perform worse than our fine-tuned LLaMA2 QLoRA model, even when prompted using a retrieval-based few-shot in-context learning method.
@@
Figure 2: Communities that consider themselves higher quality (a), more trustworthy (b), more engaged (c), more inclusive (d), and more safe (e) all use more positive and less negative sentiment to describe their moderators. Here, communities are grouped into quartiles based on their community members’ self-reported perceptions of the current state of the community. This effect is most pronounced for communities’ self-reported trustworthiness (b), with the top-25% most trustworthy communities using 34% more positive and 22% less negative language to describe their mods. Communities that rate themselves as feeling smaller (f) have a more positive perception of their mods.
@@
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4&4: The Amount of Mod Discourse for a given community and a given time period is the fraction of all posts and comments within that community made during that time period which are discussing mods, regardless of their sentiment. The Fraction of Mod Discourse with positive/negative sentiment is the fraction of all posts and comments discussing mods in a given community that have positive/negative sentiment. For many of our analyses, the time period used is the 18-month period from January 2020 through June 2021. To ensure a meaningful amount of data, we exclude communities that did not have at least one positive, neutral, and negative post or comment discussing moderators. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. These filtering steps leave us with a group of 5,282 subreddits used in downstream analyses. For some analyses, we also compute the Impact of Particular Mod’s Appointment by computing the fraction of mod discourse for the 8-week period immediately preceding and following a new moderator’s appointment to a community, and then taking the difference before vs. after they were appointed.
3.4 Ethical Considerations
We believe this work will have a positive broader impact by informing better moderation practices in online communities, as well as providing researchers with better tools to study community members’ perceptions of moderators. As we only make use of public data, we believe our work has minimal risk to participants’ privacy. As research has shown that some online community users are uncomfortable with their data being used for research, even when posted publicly (Fiesler and Proferes 2018), we take further steps to reduce potential harms and misuse potential of our mod discourse dataset: we do not publish usernames or identifiable information, only predicted sentiment with regards to the moderators. We publish moderator timelines, including moderators’ usernames, however these usernames are already publicly listed on communities’ ‘about’ pages. Upon publication of this dataset, we will provide affordances for users to have their data removed at their request. Additional dataset details, including FAIR principles, are in Appendix C. We comply with relevant licenses for NLP models we use or modify. This study was approved by the University of Washington IRB under ID number STUDY00011457.
4 How are moderators of different communities perceived differently by their communities?

### v2

@@ Abstract.
Millions of online communities are governed by volunteer moderators, who shape their communities by setting and enforcing rules, recruiting additional moderators, and participating in the community themselves. These moderators must regularly make decisions about how to govern, yet measuring the ‘success’ of governance is complex and nuanced, making it challenging to determine what governance strategies are most successful. Furthermore, prior work has shown that communities have differing values, suggesting that ‘one-size-fits-all’ approaches to governance are unlikely to serve all communities well. In this work, we assess governance practices on reddit by classifying the sentiment of community members’ public discussion of their own moderators. We label 1.89 million posts and comments made on reddit over an 18 month period. We relate these perceptions to characteristics of community governance, and to different actions that community moderators take. We identify types of communities where moderators are perceived particularly positively and negatively, and highlight promising strategies for moderator teams. Amongst other findings, we show that strict rule enforcement is linked to more favorable perceptions of moderators of communities dedicated to certain topics, such as news communities, than others. We investigate what kinds of moderators are associated with improved community perceptions upon their addition to a mod team, and find that moderators who are active community members before and during their mod tenures are seen more favorably. We make all our models, datasets, and code public.
1. Introduction
@@
Millions of online communities rely on volunteer moderators to carry out governance. These moderators perform essential services that shape their communities, including setting and enforcing rules (Fiesler et al., 2018; Chandrasekharan et al., 2018; Seering et al., 2019), communicating their actions to the community (Jhaver et al., 2019; Zhang et al., 2020), and recruiting new moderators. Some form of governance provided by moderators is essential to the healthy functioning of almost all online communities (Kraut and Resnick, 2012; Matias, 2019a).
@@
Mod teams must determine how many moderators to have, how to recruit new mods, what rules to set, and answer many more questions. With so many decisions to make, it’s challenging to determine what governance practices are most effective at ensuring high quality outcomes for online communities. There are so many different communities, with different sizes and topics, that there is no ‘one-size-fits-all’ solution (Weld et al., 2024, 2022). However, the variety of existing communities presents an opportunity: if we can develop a method to assess the success of communities’ governance practices, we can leverage the natural diversity of these practices to identify the most promising strategies for moderators. Doing so is challenging. Survey methods are effective, but expensive to deploy and cannot enable the use of historic data. Notions of ‘success’ are multifaceted, and while classifiers have been used to detect specific harms such as misinformation (Weld et al., 2021) or specific aspects of governance such as rules enforcement (Chandrasekharan et al., 2018), current methods are unable to quantify broader notions of successful governance.
@@
In this work, we measure online community governance by examining how community members themselves perceive their moderators. We develop a method to classify how community members discuss their moderators, publicly, within their communities (§3). We use this method to gather and characterize community members’ perceptions of moderators at a massive scale, enabling the largest study-to-date of governance practices that we are aware of. We label 1.89 million posts and comments from 8,477 unique subreddits across an 18-month period from January 2020-June 2021, and relate these data to different kinds of online communities, and to different actions that community mod teams can take.
@@
We identify actionable strategies for community moderators, using IPTW and DID causal inference methods to control for confounding factors including the topic and size of communities (§5). We show that communities with fewer than 5 daily posts+comments per moderator use 2.5×2.5$\times$2.5 × more positive language to describe their mods as communities with 20×20$\times$20 × more posts+comments per mod. Yet our findings do not suggest that more strict moderation improves perception of moderators; in fact, for most types of communities, a 3 percentage point increase in removed content is associated with a 9pp increase in negative language used to describe mods, although news-sharing communities are a notable exception to this trend (§6.1). We measure the impact of adding different types of moderators to a community, and find that moderators who are also active community members before and during their tenure as moderators are most strongly associated with improved perceptions of mods (§6.2).
@@
We discuss the limitations of our methods, and identify key areas for future work (§7). We make our models, datasets, and code public [^1] Contact the first author for more information. to enable further research on this important topic.
2. Related Work
@@
Measuring Community Governance. Comprehensively measuring community governance is challenging, as data access to moderator actions is often restricted, except for some types of communities, such as Wikipedia (Panciera et al., 2009) and some gaming servers (Frey and Sumner, 2019). On reddit and most popular social media platforms, however access is difficult to obtain, and, and requires separate permission from the mods of every community to be studied (Li et al., 2022a, b). In contrast, our method works for every community whose content is public.
@@
Because of these challenges, many researchers turn to surveys to study governance, sometimes qualitatively (Matias, 2019a), and sometimes at a larger scale, for instance to quantify harassment of mods (Almerekhi et al., 2020) or moderator recruiting practices (Seering and Kairam, 2022).
@@
Governance Through the Lens of Rules. Rules are often used as a lens to study governance. Large-scale analyses of rules have been used to characterize governance at the platform level on reddit (Reddy and Chandrasekharan, 2023; Fiesler et al., 2018) and on Wikipedia (Hwang and Shaw, 2022), however it is challenging to infer much about governance in specific communities, as rules constitute only a small portion of governance activity. Surveys of community members’ attitudes towards rules have been used to understand members’ attitudes towards governance more broadly (Koshy et al., 2023), however this method is challenging to scale beyond a single community. Our method can be applied to thousands of communities.
@@
Moderation Strategies on Reddit. Rules and their enforcement can also be evaluated in terms of their impact on the community as a whole (Jhaver et al., 2019; Srinivasan et al., 2019; Jhaver et al., 2023b) and their embodiment of communitywide norms (Chandrasekharan et al., 2018). Beyond rules and their enforcement, researchers have also studied other moderator strategies on reddit, including platform level decisions such as community bans and quarantines (Chandrasekharan et al., 2020; Ribeiro et al., 2020), and community level actions such as user bans (Thomas et al., 2021) and stickied posts (Matias, 2019b). Our work quantifies many of these strategies, including content removal, mod interactions with the community, and associates them with positive and negative moderation discourse.
@@
Measuring Outcomes in Online Communities. Researchers have studied the range of values that community members hold for their communities (Weld et al., 2024), yet it is challenging to accurately predict these values automatically (Weld et al., 2022). Large-scale embeddings can be used to understand community culture (Waller and Anderson, 2020), yet not governance directly. Longitudinal work has examined how communities fare with massive growth (Lin et al., 2017) and the lifecycles of their members(Danescu-Niculescu-Mizil et al., 2013), while some research focuses on smaller scale outcomes at the conversation level (Bao et al., 2021). Our work complements this literature with a large scale method for studying governance-specific outcomes.
3. Methods
@@
Reddit is a popular social media platform that is frequently studied in the computational social sciences (Proferes et al., 2021). Reddit is composed thousands of communities, known as subreddits, each with its own community norms, rules, membership, and moderators.
@@
In this work we make extensive use of publicly available reddit data from the Pushshift (Baumgartner et al., 2020). From these data, we detect and classify posts and comments which discuss moderators to produce our dataset of mod discourse (§3.3). We collect supplementary data about moderators (§3.1) and communities (§3.2).
3.1. Computing Moderator Timelines
@@
An understanding of who moderates which subreddits when is critical to our analyses. We reconstruct timelines of the 10,000 largest subreddit’s moderators using the Wayback Machine, a web archiving service provided by the Internet Archive (Archive, 2024). We collect archived snapshots of every subreddit’s moderator info page going back to 2010, totalling 30,302 snapshots, which we combine with 10,000 present day snapshots collected directly from reddit. Each snapshot includes a list of moderators’ usernames, ordered by seniority. As the Wayback Machine snapshots higher traffic pages more frequently, we have more higher temporal resolution for larger and more popular subreddits. However, as each moderator info page contains the exact timestamp that each moderator was added, we have second-level precision for new moderator appointments. After a final data-cleaning step to remove known bots, the resulting dataset of moderator timelines consists of 89,237 moderator tenures across 9,502 subreddits.
3.2. Community Topic and Health Measures
Understanding a community’s topic and health is critical for understanding that community’s perceptions of their moderators. We used a few-shot GPT-4-based classifier to classify the 10,000 largest subreddits, based on their names, into six topical categories taken from existing work: Discussion communities, hobby communities, meme communities, news communities, and video/picture-sharing communities (Weld et al., 2022). We used the manually-labeled dataset of 123 subreddits from Weld et al. (2022) for our few-shot examples, as well as to evaluate the performance of the classifier, which has 86.1% accuracy on the test set, with a macro-average $F_{1}$ score of 0.858. To understand communities’ health, we leverage a recent survey of members of 2,151 different subreddits. Survey participants were asked to rate the current state of their community on an 11-point Likert scale with regards to nine aspects of community health such as the quality of content, the trustworthiness of the community, and the safety of the community (Weld et al., 2022). We average across all responses for each subreddit to compute an overall subreddit score for each value. Our analyses of differences between communities with different topics and health aspects are in §4.
3.3. Detecting and Classifying Community Members’ Perceptions of Moderators
In this work, we quantify community members’ publicly stated perceptions of the quality of moderation by automatically detecting and classifying posts and comments which discuss moderators. Our detection and classification pipeline consists of three stages: (1) a prefilter step which uses regular expressions to identify content written by non-moderators which include the words ‘mod(s),’ ‘moderator(s);’ (2) a detection step, which detects content discussing moderators (differentiating them from those which make other use of ‘mod’, e.g., ‘video game mods’); and (3) a classification step, which classifies the sentiment of this content into positive, neutral, negative, and exclude classes. We only include content written by non-moderators as the goal of our work us to examine how non-moderators perceive their mod teams, not on how mods discuss themselves in public. We apply our pipeline, which we make public, to all reddit posts and comments made from Jan. 2020 to June 2021.
Figure 1. Determining the sentiment with regards to the moderators of comments can be very challenging.
Detection Step Details (Step 2). To differentiate posts and comments which discuss moderators from those which use ‘mod(s),’ we fine-tuned a RoBERTa-based binary classifier (Liu et al., 2019) using a manually-labeled dataset of 1,155 posts and comments, randomly sampled from the entire reddit dataset and further divided into a training set of 655 and a test set of 500 posts and comments. RoBERTa was chosen for its high-performance, and its relative ease of training and minimal compute requirements. To improve performance and provide additional context to the model, we input not only the body of comments (or title and selftext of posts), but also the name of the subreddit the comment/post was made in, as well as the parent comment that the comment being classified was in reply to, when applicable. Our fine-tuned detector model has a precision of 0.82 and a recall of 0.94 on the test set, with an $F_{1}$ score of 0.88.
Classification Step Details (Step 3). The classification step classifies posts and comments based on their sentiment with regards to the community moderators. A comment with an overall-negative sentiment may have a positive sentiment with regards to the moderators, and vice versa (Figure 1). To label posts and comments for the classification task, two annotators worked together to iteratively refine a codebook (Supplementary Materials A), then independently labeled 200 randomly sampled posts/comments. The annotators had ‘almost perfect’ inter-annotator reliability (0.85 Fleiss’ kappa) (Landis and Koch, 1977). To produce a ‘gold standard’ test set, the same two annotators independently labeled a random sample of 500 posts/comments, then discussed their disagreements until consensus was reached. A single annotator then labeled an additional 734 posts/comments for use as a training set, of which 484 were randomly sampled, and 250 were sampled based on their proximity to the decision boundary of a simpler RoBERTa model trained for the classification task.
@@ 3.3. Detecting and Classifying Community Members’ Perceptions of Moderators
Table 1. Our Classification Step model, a LLaMA 2 model fine tuned with QLoRA (Touvron et al., 2023; Dettmers et al., 2023), exceeds the performance of a retrieval-based few-shot classifier using GPT-4. Our model is also more scalable (it can be deployed locally), more affordable, and more reliable (it is not subject to prompt filtering), than GPT-4. This table compares $F_{1}$ scores for expert human labelers, retrieval-based few-shot GPT-4, and our model, alongside an empirical class distribution random baseline and a VADER-based classifier (Hutto and Gilbert, 2014).
@@
As the purpose of this method is to identify perceived moderation quality, we must use care to identify the moderators that posts/comments are discussing. On reddit, community members occasionally discuss moderators of other communities, e.g., a member of /r/gaming praising the moderators of a specific Discord server, or a member of /r/nfl complaining about the moderators of /r/seahawks. To ensure correct attribution, we decided to limit our analyses to community members talking about their own communities’ moderators. For this, we specifically trained our model to identify posts and comments which discuss the moderators of other communities, along with content that erroneously passed the detection step. We trained on this ‘exclude’ class in addition to those used for downstream analyses: positive, neutral, and negative.
@@
Figure 2. Communities that consider themselves higher quality (a), more trustworthy (b), more engaged (c), more inclusive (d), and more safe (e) all use more positive and less negative sentiment to describe their moderators. Here, communities are grouped into quartiles based on their community members’ self-reported perceptions of the current state of the community. This effect is most pronounced for communities’ self-reported trustworthiness (b), with the top-25% most trustworthy communities using 34% more positive and 22% less negative language to describe their mods. Communities that rate themselves as feeling smaller (f) have a more positive perception of their mods. In this and all other figures, points represent mean estimates alongside bootstrapped 95% confidence intervals.
To further enhance the performance of our model, we performed data augmentation using a retrieval-based few-shot classifier built with GPT-4 (OpenAI et al., 2023). We used this classifier to label an additional sample of 10,000 posts and comments, which were combined with our manually-labeled training set to fine-tune our final classification model, a 13-billion parameter LLaMA 2 model (Touvron et al., 2023) fine-tuned with QLoRA (Dettmers et al., 2023). LLaMA 2 and QLoRA were selected for their very high performance while still being feasible to fine-tune and deploy on a massive dataset. Our model was fine-tuned on an internal university HPC cluster with 2×2$\times$2 × NVIDIA a40 GPUs, which took $\approx 13$ hours. Our final model exceeds the performance of GPT-4 on all classes (Table 1). Applying the finalized Classification Step to the results from the Detection Step left us with a labeled dataset from 8,477 communities of 196,231 posts and 1,694,551 comments which discuss moderators: 175,296 with positive sentiment, 968,235 with neutral sentiment, and 747,251 with negative sentiment, and we make this dataset public.
Accurately classifying the sentiment with regards to the moderators of a post or comment is an extremely challenging task that is often heavily reliant on context and background knowledge of the community (Figure 1). As such, off-the-shelf sentiment classifiers, such as VADER, perform very poorly (Table 1). Even general purpose LLMs such as GPT-4, which obtain state-of-the-art results on standard sentiment analysis benchmarks (Kheiri and Karimi, 2023), perform worse than our fine-tuned LLaMA2 QLoRA model, even when prompted using a retrieval-based few-shot in-context learning method.
@@
Interpreting Mod Discourse. Finally, to enable downstream analyses, we define several aggregate values for each community, used in §4 & §6: The Amount of Mod Discourse for a given community is the fraction of all posts and comments discussing mods in that community, regardless of their sentiment. The Composition of Mod Discourse with positive/negative sentiment is the percent of all posts and comments discussing mods in a given community that have positive/negative sentiment. To ensure a meaningful amount of data, we exclude communities that did not have at least one positive, neutral, and negative post or comment discussing moderators. Finally, we exclude nine communities explicitly devoted to discussing moderators, as it is infeasible to differentiate discussion of those communities’ moderators from other moderators. This filtering leaves us with 5,282 subreddits used in downstream analyses.
4. How are moderators of different communities perceived differently by their communities?